OpenStudy (anonymous):

Find the area of a square if the perimeter is equal to 38 in. A. 76 in2 B. 9.5 in2 C. 38 in2 D. 90.25 in2

OpenStudy (anonymous):

\[38 = 4x\]\[x * x = area^2\]

OpenStudy (zzr0ck3r):

square has all equal sides so perimeter is 2x+2x=4x=38 what is x? area of square = x*x = x^2 so what is area?

OpenStudy (anonymous):

i am going to solve that @zzr0ck3r

OpenStudy (zzr0ck3r):

x*x = area*unites^2 not area^2 @hobbs978

OpenStudy (zzr0ck3r):

units*

OpenStudy (anonymous):

yea thanks for that clearing that up*

OpenStudy (zzr0ck3r):

@Ambbiiee 4x=38 what is x?

OpenStudy (anonymous):

x=9.5 @zzr0ck3r